# Physical simulation of Marsupial UAV-UGV Systems Connected by a Hanging Tether using Gazebo

### v1

@@ Abstract
This paper presents a ROS 2-based simulator framework for tethered UAV-UGV marsupial systems in Gazebo. The framework models interactions among a UAV, a UGV, and a winch with dynamically adjustable length and slack of the tether. It supports both manual control and automated trajectory tracking, with the winch adjusting the length of the tether based on the relative distance between the robots. The simulator’s performance is demonstrated through experiments, including comparisons with real-world data, showcasing its capability to simulate tethered robotic systems. The framework offers a flexible tool for researchers exploring tethered robot dynamics. The source code of the simulator is publicly available for the research community.
I Introduction
@@
Marsupial systems integrating unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs) are emerging as innovative solutions for a broad spectrum of complex applications. These systems offer the unique capability of operating on challenging and expansive terrain. In the case of tethered configurations, the UAV’s operational time can be significantly extended through continuous power supply via a tether connected to the UGV, effectively overcoming the limited flight duration that typically restricts small and medium-sized UAVs to a few tens of minutes. This enhancement considerably increases the mission range and operational flexibility. Prominent applications of this technology include search and rescue operations, critical infrastructure inspections, and advanced military missions. The UAV’s ability to monitor and coordinate from the air while the UGV performs specific tasks on the ground enables effective integration of aerial and terrestrial capabilities, thus optimizing the UAV’s operational duration.
@@
Developing algorithms to coordinate these marsupial systems presents significant challenges. Designing controls that precisely synchronize the actions of UAVs and UGVs requires facing challenges related to communication, synchronization, and data fusion. The complexity increases when considering the tethers connecting these vehicles, especially in modeling and controlling the dynamics of hanging or partially taut tethers, a key aspect not yet considered by existing simulators in the state-of-the-art.
@@
Figure 1: Example of the proposed Gazebo simulator deployed in a theatre model.
@@
The authors of this paper have previously tackled the problem of trajectory planning for a marsupial system where a UAV and UGV are connected by an adjustable hanging tether, as presented in [1]. The paper introduces a solution based on the Rapidly-exploring Random Trees (RRT)* and nonlinear optimization to compute collision-free paths for both the UAV and UGV while considering the tether’s dynamics and environmental obstacles. Such research highlighted the need for accurate simulation of tethered UAV-UGV, and how crucial it would be for overcoming the practical difficulties associated with real-world testing.
@@
This paper proposes a physics simulator capable of modelling the dynamics of a hanging tether of adjustable length connecting a UAV to an UGV, enabling researchers and developers to test and optimize their algorithms in a controlled environment prior to physical implementation. This not only would minimize the costs and risks associated with physical testing, but also would accelerate the development process. This work builds on [1] by leveraging the trajectories generated by this algorithm to validate the usefulness and robustness of our ROS 2-based simulator framework. Specifically, we replicated the same trajectories from their real-world experiments (Fig. 1), demonstrating how our simulator handles hanging-tether dynamics in real-time scenarios.
@@
The paper is structured as follows: Section II reviews the existing literature on marsupial systems and tether modelling. Section III describes the simulator architecture, the models used, and the characteristics of the simulation environment in Gazebo. The validation of the simulator is presented in Section IV, there, the validation scenarios, metrics, and the data obtained from these evaluations are detailed. Finally, Section VI presents the conclusions and future work, summarizing the findings, and discusses potential directions for future research.
II Related Work
@@
Marsupial robotic systems, where a UAV is deployed from a UGV, have been the subject of various studies in recent years. These systems offer unique advantages for applications in complex environments, where the ground and aerial capabilities can be leveraged for increased versatility. A prominent example is the work presented in [1], where the authors presented a path and trajectory planning solution specifically for UAV-UGV systems connected by an adjustable tether. Their approach stands out by allowing the tether to remain slack, offering the UAV greater flexibility to navigate in environments with obstacles, and thus overcoming some of the limitations found in prior works that assumed a taut tether [2].
@@
However, the development and validation of a tethered system is a complex and risky task, involving two robots with a physical link that might produce significant perturbations in the UAV control subsystem. Realistic simulators stand as a great tool to speed up such development, but simulating tethered systems, especially when considering hanging-tether dynamics, also presents several challenges. Early works on simulators, such as [5], focused on accurately representing the mechanical properties of tethers using finite element models. This approach, while useful for studying the structural behavior of tethers under different loads, lacks the integration with the kinematic control of the UAV-UGV systems, which is essential for real-world applications.
@@
Simulations of tethered systems are crucial for evaluating algorithms before deployment in real-world settings. The ability to test various control strategies, such as the dynamic tether adjustment proposed in [9], allows researchers to fine-tune their solutions without the risks associated with physical experimentation. In our case, the presented ROS 2-based simulator framework offers a unique capability by focusing on the real-time dynamics of the entire UAV-UGV system. It incorporates both manual and automated trajectory control, with dynamic adjustments of the tether’s length and slack, enabling a flexible and comprehensive testing environment. Overall, advances in Gazebo-based simulators, combined with more complex tether dynamics models, are paving the way for more accurate and efficient marsupial robotic systems.
III Gazebo Simulator
@@
  • Controllers: Each robot, the UAV and the UGV, has its own controller that receives destination points from the trajectory tracking module and executes the required movements. The tether control is embedded in the UGV. They have been tuned to closely resemble the behaviour of the robots used in the real-world experiments (see bellow), but they can be adapted to follow different dynamics. This modular design allows researchers to replace default components with their own implementations, facilitating the testing and evaluation of various control strategies within the same simulation environment.
@@ III-A Models
TABLE I: Default tether parameters
@@
The tether model used in this simulator is adapted from an open-source repository [6], specifically designed for airborne systems in Gazebo. The tether is made up of multiple connected elements, each defined by parameters such as length, radius, mass, and joint stiffness, enabling realistic dynamic behaviours. The default parameters used in the system validation can be found in Table I. These parameters, including damping and friction, can be adjusted to suit the needs of the simulation. The model is initialized with a fully retracted configuration around the winch and dynamically adjusts its length during operation. This ensures that the tether behaves realistically as it unwinds or winds based on the relative movements of the UAV and UGV. The tether’s flexibility and ability to adapt its physical properties, such as stiffness and damping, make it essential for simulating complex tethered systems. Additionally, its attachment points to both the UAV and winch are critical for maintaining consistent slack throughout the UAV-UGV coordination.
IV Simulator Evaluation
IV-A Metrics Description
In order to create a comprehensive and scalable evaluation system, a wide range of simulation data has been collected. This dataset includes information about the velocity and position of both the UGV (along with its associated winch) and the UAV, as well as the length of the tether. The data has been chosen to cover the key aspects of the system’s dynamics, focusing on movement and tether behavior. The specific metrics derived from this data are summarized in Table II.
@@ IV-A Metrics Description
These metrics provide insights into how the UGV and UAV navigate the environment, as well as the behavior of the tether as it is released and retracted.
The primary objective of this evaluation is to assess the feasibility of the simulator as a tool for comparing trajectory tracking algorithms. A key aspect of this is demonstrating the plausibility of the physical dynamics simulated, particularly those governing the behavior of the tether. The metrics presented here are crucial for verifying that the tether’s behavior is consistent with the expected physical properties, such as rigidity, elasticity, and the interaction with both the UGV and UAV during operation.
TABLE II: Simulation Metrics and Graphs
@@
IV-B Output Metrics Data
@@
V Validation Experiments
To evaluate the performance of the simulator and validate its capabilities, we designed a series of scenarios that assess various aspects of the marsupial UAV-UGV system, including stability, tether dynamics, winch operation, and coordination between aerial and ground vehicles under different circumstances. The simulations were performed on a laptop with 32 GB of RAM, a 13th Gen Intel® Core™ i7-13620H ×16 processor, and an NVIDIA GeForce RTX 4060 Laptop GPU. The scenarios are divided into three categories: Tether Model Evaluation, aimed at assessing the accuracy of the simulated tether against a real catenary curve; Simulated Scenarios, focused on testing and verifying the system’s general functionalities; and Real Scenarios that replicate real-world experiments to validate the simulator in practical applications.
An illustrative overview of the Simulated Scenarios is provided in Fig. 4, with each subfigure corresponding to the respective scenarios described below.
V-A Tether Model Evaluation
@@
We simulated the tether using three different element lengths: 0.05 m, 0.10 m, 0.15 m, and 0.20 m. For each simulation, we recorded the length of each tether element, and calculated the averaged error between the positions of the simulated tether elements and those of the theoretical catenary curve under the same conditions. This error measurement provides insights into how the granularity of the tether model influences the simulation’s ability to accurately replicate real-world tether dynamics.
TABLE III: Error between the simulated tether and the theoretical catenary at 5 m, 10 m, 15 m, and overall mean for different element lengths
Element length (m) $err_{5}$ (%) $err_{10}$ (%) $err_{15}$ (%) mean (%) real time factor (%) 0.05 0.942 0.268 0.269 0.493 25 0.10 0.332 0.214 0.232 0.259 49 0.15 0.675 0.323 0.292 0.430 70 0.20 1.600 0.490 0.552 0.881 82
As shown in Table III, the results indicate that the performance is realistic across all tested cases, with averaged errors below 1% of the tether length. We can see how, in general, as the length of the element increases, the simulation error with respect to the theoretical model also increases, as expected. However, this effect does not hold when the element length is 0.05 m. We think this behaviour is related with the increase of the simulation complexity, setting Gazebo’s solver close to its limits. Additionally, we observed that the real-time factor of the simulation improves as the element length increases, as shown in the last column of the table. This is due to the reduced computational load when simulating fewer (longer) elements. Notice that the winch model used in our simulation does not support tether element lengths greater than 0.2 m due to mechanical constraints. Despite varying the element lengths up to this limit, the precision of the simulation was not affected significantly. This suggests that even with larger element lengths within the supported range, the simulation can maintain a high degree of accuracy in approximating the real catenary curve.
V-B Simulated Scenarios
@@
  • Scenario 1 (Fig.  4a): Vertical Stability Assessment. In this scenario, the UGV remains stationary while the UAV performs ten consecutive ascents and descents. The objective is to evaluate the winch’s ability to smoothly release and retract the tether in response to vertical movements of the UAV. This scenario focuses on the stability of the UAV during tether length adjustments and the responsiveness of the winch control system.
@@
According to Table IV, the UGV traveled significant distances in Scenarios 2 and 3, covering 49.99 meters and 50.88 meters, respectively. The UGV consistently adhered to the designated paths, exhibiting smooth and stable movements that reflect reliable ground dynamics.
@@
Figure 5: Visualization of the position of the UAV and UGV on each axis with respect to the reference in Scenario 3 (Opposite Direction Coordination).
Fig. 6 depicts the tether length adjustments throughout the simulated scenario. The tether length is dynamically adjusted by the winch system to be 5% greater than the relative distance between the UAV and UGV. This strategy ensures that the tether maintains appropriate slack, preventing excessive tightness that could adversely affect the UAV’s stability. The tether length closely follows the target length, indicating the winch’s responsiveness and the effectiveness of the tether management algorithms in real-time operation.
V-C Real Scenarios
Figure 7: 3D representation in Gazebo of the theatre where the experiment was carried out.
Building upon the insights gained from the simulated scenarios, we proceeded to more complex tests by replicating real-world experiments. Scenarios 4 and 5 focus on replicating a complex trajectory within a confined environment, specifically designed to validate the simulator against real-world scenarios.
  • Scenario 4: Dynamic Tether Length Trajectory Replication. This scenario replicates the trajectory described in [1], simulating a movement within a theatre environment as can be seen in Fig. 7. The tether length is dynamically calculated based on the relative distance between the UAV and UGV throughout the mission. The purpose is to validate the simulator’s accuracy in modeling real-world scenarios where the tether length adjusts in real-time.
  • Scenario 5: Predefined Tether Length Trajectory Replication. Similar to scenario 4, this test replicates the trajectory from [1]. However, in this case, the tether length is predefined and provided as part of the trajectory data, rather than being dynamically calculated. This allows for a direct comparison between the simulator’s performance using dynamic versus static tether length control, highlighting the effects on system behavior and synchronization.
The results indicate that the simulated trajectories closely align with those observed in the actual experiments. The UAV and UGV successfully navigate the complex environment, negotiating obstacles and coordinating movements effectively. The simulator accurately reproduces the dynamics of both robots, including the interaction with the tether, which is critical in confined spaces where precise control is necessary.
Figure 8 shows the 3D trajectories of the UAV and UGV during scenario 5. Comparative analysis reveals that the deviations between the simulated and real-world trajectories are marginal. The UAV maintains stability despite the complex tether dynamics, and the UGV follows the planned path with high fidelity. The tether management system effectively adjusts the tether length in real-time, preventing excessive slack or tightness that could impede the robots’ movements or compromise safety.
TABLE IV: Scenario results.
Scenario Sim. time Number of targets Distance UAV Distance UGV Tether released Tether collected
@@ V-C Real Scenarios
4 628 100 33.92 16.45 17.91 7.54
5 609 100 33.96 16.41 19.57 8.40
VI Conclusions and Future Work

### v2

# Physical simulation of Marsupial UAV-UGV Systems Connected by a Variable-Length Hanging Tether

@@ Abstract
This paper presents a simulation framework able of modeling the dynamics of a hanging tether with adjustable length, connecting a UAV to a UGV. The model incorporates the interaction between the UAV, UGV, and a winch, allowing for dynamic tether adjustments based on the relative motion of the robots. The accuracy and reliability of the simulator are assessed through extensive experiments, including comparisons with real-world experiment, to evaluate its ability to reproduce the complex tether dynamics observed in physical deployments. The results demonstrate that the simulation closely aligns with real-world behavior, particularly in constrained environments where tether effects are significant. This work provides a validated tool for studying tethered robotic systems, offering valuable insights into their motion dynamics and control strategies.
keywords:
Simulation and Animation, Aerial Systems: Applications, Autonomous Vehicle Navigation
footnotetext: This work was partially supported by the grants: 1) INSERTION PID2021-127648OB-C31, and 2) NORDIC TED2021-132476B-I00 projects, funded by MCIN/AEI/ 10.13039/501100011033 and the “European Union NextGenerationEU/PRTR”
1 Introduction
@@
Marsupial systems, where unmanned aerial vehicles (UAVs) are deployed from unmanned ground vehicles (UGVs) to combine their complementary capabilities, are emerging as innovative solutions for a broad spectrum of complex applications. These systems offer the unique capability of operating on challenging and expansive terrain. In the case of tethered configurations, the UAV’s operational time can be significantly extended through continuous power supply via a tether connected to the UGV, effectively overcoming the limited flight duration that typically restricts small and medium-sized UAVs to a few tens of minutes. This enhancement considerably increases the mission range and operational flexibility. Prominent applications of this technology include search and rescue operations, critical infrastructure inspections, and advanced military missions. The UAV’s ability to monitor and coordinate from the air while the UGV performs specific tasks on the ground enables effective integration of aerial and terrestrial capabilities, thus optimizing the UAV’s operational duration.
@@
However, tethered configurations inherently limits maneuverability. To overcome this constraint, advanced planning algorithms and appropiate simulation framework are required. Developing these algorithms to coordinate these marsupial systems presents significant challenges. Designing controls that precisely synchronize the actions of UAVs and UGVs requires facing challenges related to communication, synchronization, and data fusion. The complexity increases when considering the tethers connecting these vehicles, especially in modeling and controlling the dynamics of hanging or partially taut tethers, a key aspect not yet considered by existing simulators in the state-of-the-art.
@@
Figure 1: Example of the proposed marsupial system simulator deployed in a theatre model using Gazebo.
@@
The authors of this paper have previously tackled the problem of trajectory planning for a marsupial system where a UAV and UGV are connected by an adjustable hanging tether, as presented in [1]. The paper introduces a solution based on the Rapidly-exploring Random Trees (RRT) and nonlinear optimization to compute collision-free paths for both the UAV and UGV while considering the tether’s dynamics and environmental obstacles. While this approach was effective for planning, it lacked a detailed physics-based simulation of tether behaviour under various real-world conditions. Consequently, such research highlighted the need for accurate simulation of tethered UAV-UGV, and how crucial it would be for overcoming the practical difficulties associated with real-world testing.
@@
This paper proposes and studies the viability of a physics simulator able of modelling the dynamics of a hanging tether of adjustable length connecting a UAV to an UGV, enabling researchers and developers to test and optimize their algorithms in a controlled environment prior to physical implementation. This not only would minimize the costs and risks associated with physical testing, but also would accelerate the development process. This simulation is validated using the work on [1], by leveraging the trajectories generated by the algorithm presented there to check the usefulness and robustness of our ROS 2-based simulator framework. Specifically, we replicated the same trajectories from their real-world experiments (Fig. 1), demonstrating how our simulator handles hanging-tether dynamics in real scenarios.
@@
The paper is structured as follows: Section 2 reviews the existing literature on marsupial systems and tether modelling. Section 3 describes the simulator architecture, the models used, and the characteristics of the simulation environment in Gazebo. The metrics of the simulator are presented in Section 4, there, the metrics and the data obtained are detailed. Section 5 evaluates the simulator’s capabilities through experiments on tether model accuracy, system behavior, and computational performance, including a comparison with real-world data. Finally, Section 6 presents the conclusions and future work, summarizing the findings, and discusses potential directions for future research.
2 Related Work
@@
Marsupial robotic systems, where a UAV is deployed from a UGV, have been the subject of various studies in recent years. These systems offer unique advantages for applications in complex environments, where the ground and aerial capabilities can be leveraged for increased versatility. Early works, such as [2], assumed a taut tether, which restricted the UAV’s maneuverability in obstacle-rich environments. In contrast, [1] presented a path and trajectory planning solution specifically for UAV-UGV systems connected by an adjustable tether. Their approach allows the tether to remain slack, providing the UAV with greater flexibility and overcoming some of the limitations found in prior studies.
@@
However, the development and validation of a tethered system is a complex and risky task, involving two robots with a physical link that might produce significant perturbations in the UAV control subsystem. Realistic simulators stand as a great tool to speed up such development, but simulating tethered systems, especially when considering hanging-tether dynamics, also presents several challenges. Early works on simulators, such as [5], focused on accurately representing the mechanical properties of tethers using finite element models. This approach, while useful for studying the structural behavior of tethers under different loads, lacks the integration with the kinematic control of the UAV-UGV systems, which is essential for real-world applications. In contrast, our approach prioritizes system-level simulation, allowing for the analysis of robot interactions, control strategies, and tether behavior in a comprehensive and operationally relevant manner.
@@
Simulations of tethered systems are crucial for evaluating algorithms before deployment in real-world settings. The ability to test various control strategies, such as the tether-aware path planning method in [9], allows researchers to fine-tune their solutions without the risks associated with physical experimentation. In our case, the presented ROS 2-based simulator framework offers a unique capability by focusing on the real-time dynamics of the entire UAV-UGV system. It incorporates both manual and automated trajectory control, with dynamic adjustments of the tether’s length and slack, enabling a flexible and comprehensive testing environment. Overall, advances in Gazebo-based simulators, combined with more complex tether dynamics models, are paving the way for more accurate and efficient marsupial robotic systems.
3 Marsupial System Simulator using Gazebo
@@
  • Controllers: Each robot, the UAV and the UGV, has its own controller that receives destination points from the trajectory tracking module and executes the required movements. The tether control is embedded in the UGV. They have been tuned to closely resemble the behaviour of the robots used in the real-world experiments (see below), but they can be adapted to follow different dynamics. This modular design allows researchers to replace default components with their own implementations, facilitating the testing and evaluation of various control strategies within the same simulation environment.
@@ 3.1 Models
Table 1: Default tether parameters
@@
The tether model used in this simulator is adapted from an open-source repository [6], specifically designed for airborne systems in Gazebo. The tether is made up of multiple connected elements, each defined by parameters such as length, radius, mass, and joint stiffness, enabling realistic dynamic behaviours. The default parameters used in the system validation can be found in Table 1. These parameters, including damping and friction, can be adjusted to suit the needs of the simulation. The model is initialized with a fully retracted configuration around the winch and dynamically adjusts its length during operation. This ensures that the tether behaves realistically as it unwinds or winds based on the relative movements of the UAV and UGV. The tether’s flexibility and ability to adapt its physical properties, such as stiffness and damping, make it essential for simulating complex tethered systems. Additionally, its attachment points to both the UAV and winch are critical for maintaining consistent slack throughout the UAV-UGV coordination.
Figure 4: Example of tether-obstacle interaction in Gazebo.
Beyond its dynamic adjustments, the tether model also interacts with the environment through Gazebo’s physics engine, which computes contact forces based on material properties and collision geometry. These interactions create local tension variations that propagate along the tether, potentially affecting UAV and UGV dynamics. As illustrated in Fig. 4, Gazebo naturally handles tether-obstacle interactions, enabling realistic responses to environmental constraints. To assess these effects, the simulator includes six scenarios with different architectural structures, enabling analysis of tether behavior under varied geometric conditions.
4 Metrics for Simulation Evaluation
In order to create a comprehensive and scalable evaluation system, a wide range of simulation data has been collected. The primary objective of this evaluation is to assess the feasibility of the simulator as a tool for comparing trajectory tracking algorithms. A key aspect of this is demonstrating the plausibility of the simulated physical dynamics, particularly those governing the behavior of the tether.
4.1 Metrics Description
This dataset includes information about the velocity and position of both the UGV (along with its associated winch) and the UAV, as well as the length of the tether. The data has been chosen to cover the key aspects of the system’s dynamics, focusing on movement and tether behavior. The specific metrics derived from this data are summarized in Table 2.
@@ 4.1 Metrics Description
Beyond the tether shape itself, these metrics also provide insights into how the UGV and UAV navigate the environment, as well as how the tether behaves as it is released and retracted during motion. Ultimately, the metrics presented here are crucial for verifying that the tether’s behavior remains consistent with expected physical properties, such as rigidity, elasticity, and its interaction with both the UGV and UAV throughout operation.
Table 2: Simulation Metrics and Graphs
@@
[…]
4.2 Computed Metrics Data
@@
5 Validation Experiments
To evaluate the performance of the simulator and validate its capabilities, we designed a series of scenarios that assess various aspects of the marsupial UAV-UGV system, including stability, tether dynamics, winch operation, and coordination between aerial and ground vehicles under different circumstances. The scenarios are divided into four categories: Tether Model Evaluation, aimed at assessing the accuracy of the simulated tether against a real catenary curve; Simulated Scenarios, focused on testing and verifying the system’s general functionalities; Real Scenarios, that replicate real-world experiments to validate the simulator in practical applications; and Performance Evaluation, dedicated to computational efficiency and real-time capabilities of the simulator under different conditions.
The documentation in our GitHub repository details how to configure new environments, robot initial poses, motion trajectories, and tether parameters, allowing researchers to adapt the framework to a wide range of experimental setups. Additionally, a summary of the simulation and real-world experiments is presented in the accompanying video [^2] https://youtu.be/ZLLDROIaHV0, providing a visual overview of the system’s capabilities. This flexibility ensures the simulator remains a valuable and continually evolving tool for marsupial UAV-UGV research.
5.1 Tether Model Evaluation
@@
We simulate the tether using four different lengths of elements: 0.05 m, 0.10 m, 0.15 m and 0.20 m. For each simulation, we recorded the length of each tether element and calculated the averaged error between the positions of the simulated tether elements and those of the theoretical catenary curve under the same conditions. This error measurement provides insights into how the granularity of the tether model influences the simulation’s ability to accurately replicate real-world tether dynamics.
Table 3: Error between the simulated tether and the theoretical catenary for tether lengths of 5 m, 10 m, 15 m, and different element lengths (mean values)
Element length (m) $err_{5}$ (%) $err_{10}$ (%) $err_{15}$ (%) mean (%)
0.05 0.942 0.268 0.269 0.493
0.10 0.332 0.214 0.232 0.259
0.15 0.675 0.323 0.292 0.430
0.20 1.600 0.490 0.552 0.881
[…]
As shown in Table 3, the results indicate that the performance is realistic across all tested cases, with averaged errors below 1% of the tether length. We can see how, in general, as the length of the element increases, the simulation error with respect to the theoretical model also increases, as expected. However, this effect does not hold when the element length is 0.05 m. We think this behaviour is related with the increase of the simulation complexity, setting Gazebo’s solver close to its limits. This is due to the reduced computational load when simulating fewer (longer) elements.
It is important to note that the region around the winch (highlighted in yellow in Fig. 3) exhibits more chaotical tether dynamics due to the high concentration of elements in a small area. This effect is inherent to the discrete nature of the simulation and cannot be directly controlled, leading to small fluctuations in the tether’s appearance near the winch. However, this is purely a visual artifact and does not impact the overall behaviour of the catenary, which remains accurately modeled and is the primary focus of our simulation. Additionally, the winch model used in our simulation does not support tether element lengths greater than 0.2 m due to mechanical constraints. Despite varying the element lengths up to this limit, the precision of the simulation was not affected significantly. This suggests that even with larger element lengths within the supported range, the simulation can maintain a high degree of accuracy in approximating the real catenary curve.
(a) Collision with a hanging tether.
(b) Collision with a taut tether.
Figure 5: Examples of tether-obstacle collisions with varying degrees of tension.
Beyond comparing the simulated tether to a theoretical catenary, we also conducted experiments to examine how it interacts with the environment. These tests were designed to confirm the simulator’s ability to capture contact dynamics, including collisions with obstacles, wrapping around structures, and adjusting the tether’s tension in response. Thus, Fig. 5(a) illustrates a scenario where the tether remains slack while suspended between two obstacles. In contrast, Fig. 5(b) shows the tether becoming entangled around both obstacles due to the UAV’s trajectory, significantly increasing tension. These interactions influence the overall system behavior, affecting both the UAV’s stability and the UGV’s traction due to the dynamic forces exerted by the tether. This behavior can be better appreciated in the accompanying video [^3] https://youtu.be/_j3KnPfFLBQ.
The ability to simulate these contact events is essential for realistic modeling of marsupial robotic systems, as tether-environment interactions can introduce significant constraints in real-world applications. Our results demonstrate that the simulator correctly captures these effects, enabling detailed analysis of how tether dynamics impact system performance under various operational conditions.
5.2 Simulated Scenarios
@@
  • Scenario 1 (Fig.  6a): Vertical Stability Assessment. In this scenario, the UGV remains stationary while the UAV performs ten consecutive ascents and descents. The objective is to evaluate the winch’s ability to smoothly release and retract the tether in response to vertical movements of the UAV. This scenario focuses on the stability of the UAV during tether length adjustments and the responsiveness of the winch control system.
@@
According to Table 4, the UGV traveled significant distances in Scenarios 2 and 3, covering 49.99 meters and 50.88 meters, respectively. The UGV consistently adhered to the designated paths, exhibiting smooth and stable movements that reflect reliable ground dynamics.
@@
Specifically, the table reports key performance metrics: the simulation time (in seconds), the number of targets reached, the total distance traveled by the UAV and UGV (in meters), and the amount of tether released and collected (in meters) throughout the experiment
Table 4: Simulation results of the real experiment scenario. Key performance metrics include simulation time (in seconds), number of targets reached, total distance traveled by the UAV and UGV (in meters), and the amount of tether released and collected (in meters) throughout the experiment.
Scenario Simulation time Number of targets Distance UAV Distance UGV Tether released Tether collected
@@
[…]
Figure 7: Visualization of the position of the UGV (left) and UAV (right) on each axis with respect to the reference in Scenario 3 (Opposite Direction Coordination).
Figure 8 depicts the tether length adjustments throughout the simulated scenario. The tether length is dynamically adjusted by the winch system to be 5% greater than the relative distance between the UAV and UGV. This strategy ensures that the tether maintains slightly slack, preventing excessive tightness that might impact the UAV’s stability. The tether length closely follows the target length, indicating the winch’s responsiveness and the effectiveness of the tether management algorithms in real-time operation.
5.3 Comparison with real experiment
Figure 9: 3D Lidar reconstruction of the theater in Gazebo where the experiment was conducted (up) and the real-world experiment used as a reference to validate the proposed simulation framework (down).
To evaluate the realism and accuracy of the proposed simulator, a real-world experiment with a marsupial robotic system was first conducted by our team (see Fig. 9), then replicated in simulation, and subsequently compared to the real experiment. This experiment, detailed in [1], was conducted with a physical tethered UAV-UGV system navigating a confined environment, a theatre. The marsupial system consisted of a UAV and a UGV connected by an adjustable hanging tether, with both the UAV and the UGV actively navigating, while the UGV also provided power and managed the tether. The experiment demonstrated the feasibility of planning and executing collision-free trajectories in a real-world setting, highlighting the challenges posed by tether dynamics and confined spaces.
Using the same parameters and trajectory plans as those employed in the real-world experiment, we replicated the setup in our simulator. The goal was to assess the simulator’s ability to model the behavior of the tether and the coordinated movements of the UAV and UGV in a highly constrained environment. The theatre environment was recreated in Gazebo using a Lidar reconstruction, including obstacles and the same initial conditions, as shown in Fig. 9.
To further evaluate the simulator’s performance, two specific scenarios were designed based on the replicated environment. In the first scenario, Real-Time Tether Adaptation (RTTA), the tether length was dynamically adjusted throughout the mission proportionally to the distance between the vehicles, adapting in real time to the relative positions of the UAV and UGV. This approach allowed for continuous adjustments to ensure proper slack and operational stability. In the second scenario, Predefined Tether Reference (PTR), the same trajectory was replicated using a pre-planned tether length for each waypoint of the trajectory, mirroring the behaviour of the real-world system, where the tether length is planned as part of the trajectory.
The comparison shown in Table 5 between these two scenarios offered valuable insights into how different tether management strategies affect system behavior and synchronization. The values for tether released and collected differ between the RTTA and PTR scenarios due to their distinct tether management strategies. Notably, the greatest discrepancies between simulation and real-world experiments occur when the UAV and UGV are further apart, as the extended tether introduces additional dynamic effects that influence the UAV’s motion. However, these differences remain minimal, and the overall trajectory alignment demonstrates the simulator’s ability to accurately capture key aspects of real-world marsupial robotic behavior. The UAV and UGV successfully navigate the complex environment, negotiating obstacles and coordinating movements effectively. This analysis underscored the versatility and fidelity of the simulator in modeling complex marsupial robotic systems.
Figure 10 shows the 3D trajectories of the UAV and UGV during PTR scenario. Comparative analysis reveals that the deviations between the simulated and real-world trajectories are marginal. The UAV maintains stability despite the complex tether dynamics, and the UGV follows the planned path with high fidelity. The tether management system effectively adjusts the tether length in real-time, preventing excessive slack or tightness that could impede the robots’ movements or compromise safety.
(a) UGV comparison
(b) UAV comparison
Table 5: Simulation results of the real experiment scenario. Key performance metrics include simulation time (in seconds), number of targets reached, total distance traveled by the UAV and UGV (in meters), and the amount of tether released and collected (in meters) throughout the experiment.
Scenario Simulation time Number of targets Distance UAV Distance UGV Tether released Tether collected
RTTA 628 100 33.92 16.45 17.91 7.54
PTR 609 100 33.96 16.41 19.57 8.40
[…]
5.4 Performance Evaluation
The computational performance of the simulator was evaluated to assess its real-time capabilities and scalability. All experiments were conducted on a laptop with 32 GB of RAM, a 13th Gen Intel Core i7-13620H (10 cores, 16 threads), and an NVIDIA GeForce RTX 4060 Laptop GPU. To analyze the impact of system components on simulation efficiency, a comparison was made with a more powerful desktop PC (64 GB RAM, 12th Gen Intel Core i9-12900F, and an NVIDIA GeForce RTX 3060 GPU). Each experiment was repeated 10 times, and since the mean deviation was negligible, the values correspond to the calculated averages
Figure 11: Performance comparison between the laptop and desktop configurations. The real-time factor (RTF) is shown for tether element counts ranging from 100 to 700.
As shown in Fig. 11, the desktop PC demonstrated superior performance. The number of tether elements directly influenced computational load. For the laptop, the real-time factor (RTF) decreased from 0.92 (100 elements) to 0.14 (700 elements), while the desktop PC maintained higher RTF values, ranging from 0.99 (100 elements) to 0.19 (700 elements), with a tether element length of 0.15 m in all cases. This divergence underscores the importance of system specifications for simulating high-fidelity tethered systems. Notably, the desktop achieved an RTF of 0.47 with 300 elements while the laptop required element counts below 200 to maintain RTF  0.50.
These results highlight the simulator’s ability to run on mid-tier hardware while showcasing the benefits of higher-end systems for demanding simulations. As the number of tether elements increases, maintaining real-time performance becomes progressively more demanding. While the simulator remains functional across a range of systems, achieving high real-time factors, which is necessary if real-time execution is required, demands sufficient computational resources. Therefore, selecting the appropriate hardware is crucial to balancing simulation accuracy and computational feasibility in real-time applications.
6 Conclusions and Future Work